Twenty pounds of lemon drops are to be mixed with cherry chews to make a mixture that will sell for $1.80 per pound. How much of the more expensive candy should be used?
Given Information:
Peppermint Patties= $1.35 price per pound
Lemon Drops= $1.70 price per pound
Licorice lumps= $1.95 price per pound
Cherry Chews= $2.00 price per pound

Answer by TimothyLamb(4379)

---
x = pounds lemon
y = pounds cherry
x = 20
---
1.7x + 2y = 1.8(x + y)
1.7x + 2y = 1.8x + 1.8y
2y - 1.8y = 1.8x - 1.7x
0.2y = 0.1x
y = (0.1/0.2)x
y = (0.1/0.2)*20
---
y = pounds cherry = 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php